# Einstein from Noise: Statistical Analysis

### v2

@@ H.3 Convergence of the real and imaginary parts of the EfN estimator
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\mathbb{E}\left(\left[\left|\mathsf{Z}_{1}[k]\right|\sin(\phi_{e,\mathsf{Z}_{1}}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{Z}_{1}[k]\right|\cos(\phi_{e,\mathsf{Z}_{1}}[k])\right]}\right)^{2}}.$ (H.36)
which is similar to (B.24). Applying (H.29) and (H.30) into (H.36) yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\lim_{d\to\infty}\frac{\mathbb{E}\left(\left[\left|\mathsf{N}_{1}[k]\right|\sin(\phi_{e,\mathsf{N}_{1}}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{N}_{1}[k]\right|\cos(\phi_{e,\mathsf{N}_{1}}[k])\right]}\right)^{2}}=C_{k}<\infty.$ (H.37)
Finally, for (5.5), under the assumption that $x$ satisfies Assumption 4.2, then by (4.6), the r.h.s. of (H.37) converges to,
$\lim_{d\to\infty}a^{2}_{d}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}=\frac{1}{2},$ (H.38)
where $a_{d}=\sqrt{2\log(d)}$. Thus, substituting (H.38) into the r.h.s. of (H.37) yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}\left[|\phi_{\hat{\mathsf{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}\right]}{1/(M\log d)}\cdot\frac{1}{1/(4|\mathsf{X}[k]|^{2})}=1,$ (H.39)
@@
Appendix I Proof of Proposition 5.4
The proof strategy for Proposition 5.2 closely follows that of the i.i.d. Gaussian case (Theorem 4.1), with appropriate modifications to handle circulant noise. The necessary assumptions and notations are introduced in Appendix 5.4. Appendix I.2 establishes the asymptotic convergence of the EfN estimator as $M\to\infty$ under circulant Gaussian noise statistics. In Appendix I.3, we show that conditioning the EfN process on a single Fourier noise coefficient results in a cyclo-stationary process with a cosine trend. Appendix I.4 extends the vanishing imaginary part result from Appendix B.1 to the setting of circulant Gaussian noise. Similarly, Appendix I.5 extends the result of Appendix B.2, showing that the real part remains strictly positive in the circulant case. Finally, Appendix I.6 combines the results of the preceding sections to complete the proof of Proposition 5.4.
I.1 Preliminaries
@@
    Fourier phases distribution. For any $k$ such that $\lambda_{k}>0$, the Fourier coefficient $\mathsf{Y}_{i}[k]$ is a zero-mean, circularly symmetric complex Gaussian random variable. Hence, the phases $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d/2}$ are i.i.d. and uniformly distributed on $[-\pi,\pi)$ and independent of the magnitude $\{|\mathsf{Y}_{i}[k]|\}_{k=0}^{d/2}$.
@@
    $\mathsf{Y}_{i}[d-k]=\overline{\mathsf{Y}_{i}[k]},\quad\text{for }1\leq k\leq d-1.$
@@
I.2 The convergence of the Einstein from Noise estimator for circulant Gaussian noise
@@
$\hat{\mathsf{X}}[k]$ $=\frac{1}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{\mathsf{N}_{i}}[k]}\,e^{j\frac{2\pi k}{d}\mathsf{\hat{R}}_{i}}$ (I.1)
$=\frac{e^{j\phi_{\mathsf{X}}[k]}}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{\mathsf{N}_{i}}[k]}\,e^{j\frac{2\pi k}{d}\mathsf{\hat{R}}_{i}}\,e^{-j\phi_{\mathsf{X}}[k]}$ (I.2)
$=\frac{e^{j\phi_{\mathsf{X}}[k]}}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{e,i}[k]},$ (I.3)
@@
$\mathsf{\hat{R}}_{i}\triangleq\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (I.4)
@@
$\phi_{e,i}[k]\triangleq\frac{2\pi k\mathsf{\hat{R}}_{i}}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k].$ (I.5)
@@
$\mathsf{S}_{i}[r]\triangleq\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (I.6)
so that $\mathsf{\hat{R}}_{i}=\operatorname*{arg\,max}_{0\leq r\leq d-1}\mathsf{S}_{i}[r]$. We note that for any $10\leq i\leq M-$ 1, the random vector $\mathsf{S}_{i}\triangleq(\mathsf{S}_{i}[0],\mathsf{S}_{i}[1],\ldots,\mathsf{S}_{i}[d-1])^{\top}$ is jointly Gaussian with zero mean and a circulant covariance matrix. Hence, $\mathsf{S}_{i}$ forms a cyclo-stationary process. Applying the strong law of large numbers (SLLN), as $M\to\infty$, we have,
$\mathsf{\hat{X}}[k]\,e^{-j\phi_{\mathsf{X}}[k]}$ $=\frac{1}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{e,i}[k]}$ (I.7)
[…]
@@
$\mu_{\mathsf{A},k}$ $\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right],$ (I.9)
$\mu_{\mathsf{B},k}$ $\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right],$ (I.10)
@@
I.3 Conditioning on the Fourier frequency noise component
@@
Lemma I.2.
Let $\mathsf{S}_{i}$ be defined as in (I.6), and denote $\mathbb{E}\left[|\mathsf{Y}_{i}[k]|^{2}\right]=\lambda_{k}>0$ for each $k\in\{0,1,\ldots,d-1\}$. Then, for every $k\in\left\{1,2,\ldots,\tfrac{d}{2}-1,\tfrac{d}{2}+1,\ldots,d-1\right\}$, the random vector $\mathsf{S}_{i}$ conditioned on $\mathsf{Y}_{i}[k]$ is Gaussian:
$\mathsf{S}_{i}|\mathsf{Y}_{i}[k]\sim\mathcal{N}(\mathsf{\mu}_{k,i},\mathsf{\Sigma}_{k,i}),$ (I.11)
@@
$\mu_{k,i}[r]$ $\triangleq\mathbb{E}\left[\mathsf{S}_{i}[r]|\mathsf{Y}_{i}[k]\right]=2\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (I.12)
for $10\leq r\leq d-$ 1, and
@@
$=\sum_{\ell=0}^{d-1}\lambda_{\ell}\cdot|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$ (I.13)
for $10\leq r,s\leq d-$ 1, where $\tilde{\mathsf{X}}_{k}$ is defined by:
$\tilde{\mathsf{X}}_{k}[\ell]\triangleq\begin{cases}0&\mathsf{if}\;\ell=k,d-k,\\ \mathsf{X}[\ell]&\mathsf{if}\;\ell=0,d/2,\\ \sqrt{2}\cdot\mathsf{X}[\ell]&\mathsf{otherwise}.\end{cases}$ (I.14)
Note that the conditional process $\mathsf{S}_{i}|\mathsf{Y}_{i}[k]$ is Gaussian because it is given by a linear transform of i.i.d. Gaussian variables. Also, since its covariance matrix is circulant and depends only on the difference between the two indices, i.e., $\mathsf{\Sigma}_{k,i}[r,s]=\sigma_{k,i}[\left|r-s\right|]$, it is cycle-stationary with a cosine trend. The eigenvalues of this circulant matrix are given by the DFT of its first row, and thus its $\ell$-th eigenvalue equals $\lambda_{\ell}\cdot|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$, for $10\leq\ell\leq d-$ 1.
Remark I.3.
@@
Proof of Lemma I.2.
We recall that if $\left\{y_{i}\right\}_{i=0}^{M-1}\sim\mathcal{N}(0,\Sigma)$, for symmetric circulant matrix $\Sigma$, then their DFT coefficients satisfy ${\{|\mathsf{Y}_{i}[k]|}\}_{k=0}^{d-1}$, and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}$ are independent and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}\sim\mathsf{Unif}[-\pi,\pi)$. By definition of $\mathsf{S}_{i}$ (I.6), we have for every $k\neq 0,d/2$,
@@
$+\sum_{\ell\neq k,d-k}{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)},$ (I.15)
where we have used the property of $\mathsf{X}[k]=\overline{\mathsf{X}[d-k]},\ \mathsf{Y}_{i}[k]=\overline{\mathsf{Y}_{i}[d-k]}$. Clearly, as $\mathbb{E}\left[\mathsf{Y}_{i}\left[\ell\right]\right]=0$, for every $10\leq\ell\leq d-$ 1, we have,
$\mathbb{E}\left[{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)}\right]=0,$ (I.16)
for every $10\leq\ell\leq d-$ 1. Combining (I.15) and (I.16) results,
$\mu_{k,i}[r]$ $=\mathbb{E}\left[\mathsf{S}_{i}[r]|\mathsf{Y}_{i}[k]\right]=2\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (I.17)
@@
$\mathsf{\Sigma}_{k,i}[r,s]$ $\triangleq\mathbb{E}\left[\left(\mathsf{S}_{i}[r]-\mathbb{E}\mathsf{S}_{i}[r]\right)\left(\mathsf{S}_{i}[s]-\mathbb{E}\mathsf{S}_{i}[s]\right)|\mathsf{Y}_{i}[k]\right].$ (I.18)
@@
$=\sum_{\ell\neq k,d-k}{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)},$ (I.19)
@@
=$\mathcal{I}=\left\{1,2,\dots k-1,k+1,\dots,d/2-1\right\},$ (I.20)
@@
$+2\cdot\sum_{\ell\in\mathcal{I}}\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right),$ (I.21)
where each one of the terms in the sum is independent. Since the terms in the sum on the r.h.s. of (I.21) are independent—that is, $\mathbb{E}\left[\mathsf{Y}_{i}\left[\ell_{1}\right]\mathsf{Y}_{i}\left[\ell_{2}\right]\right]=\mathbb{E}\left[\left|\mathsf{Y}_{i}\left[\ell_{1}\right]\right|^{2}\right]\delta_{\ell_{1},\ell_{2}}$—it follows that,
@@
$+4\cdot\mathbb{E}\left[{\sum_{\ell\in\mathcal{I}}{\left|\mathsf{X}[\ell]\right|^{2}\left|\mathsf{Y}_{i}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\cos\left(\frac{2\pi\ell s}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)}}\right].$ (I.22)
The expectation value in (I.22) is composed of the multiplications of cosines. Applying trigonometric identities, we obtain,
@@
$=\cos\left(\frac{2\pi\ell(r-s)}{d}\right)+\cos\left(\frac{2\pi\ell(r+s)}{d}+2\left(\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\right),$ (I.23)
for every $10\leq r,s\leq d-$ 1. Now, since the sequences $\left\{|\mathsf{Y}_{i}[\ell]|\right\}_{\ell=0}^{d/2}$ and $\left\{\phi_{\mathsf{Y}_{i}}[\ell]\right\}_{\ell=0}^{d/2}$ are independent random variables, with $\mathbb{E}\left[\left|\mathsf{Y}_{i}[k]\right|^{2}\right]=\lambda_{k}$ and phases $\phi_{\mathsf{Y}_{i}}[k]$ uniformly distributed over $[-\pi,\pi)$—and by applying the trigonometric identity (I.23), it follows that,

### v3

@@ F.3 Convergence of the real and imaginary parts of the EfN estimator
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\mathbb{E}\left(\left[\left|\mathsf{Z}_{1}[k]\right|\sin(\phi_{e,\mathsf{Z}_{1}}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{Z}_{1}[k]\right|\cos(\phi_{e,\mathsf{Z}_{1}}[k])\right]}\right)^{2}}.$ (F.38)
which is similar to (B.24). Applying (F.31) and (F.32) into (F.38) yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\lim_{d\to\infty}\frac{\mathbb{E}\left(\left[\left|\mathsf{N}_{1}[k]\right|\sin(\phi_{e,\mathsf{N}_{1}}[k])\right]^{2}\right)}{\left(\mathbb{E}{\left[\left|\mathsf{N}_{1}[k]\right|\cos(\phi_{e,\mathsf{N}_{1}}[k])\right]}\right)^{2}}=C_{k}<\infty.$ (F.39)
Finally, for (5.5), under the assumption that $x$ satisfies Assumption 4.2, then by (4.8), the r.h.s. of (F.39) converges to,
$\lim_{d\to\infty}a^{2}_{d}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}=\frac{1}{2},$ (F.40)
where $a_{d}=\sqrt{2\log(d)}$. Thus, substituting (F.40) into the r.h.s. of (F.39) yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\frac{\mathbb{E}\left[|\phi_{\hat{\mathsf{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}\right]}{1/(M\log d)}\cdot\frac{1}{1/(4|\mathsf{X}[k]|^{2})}=1,$ (F.41)
@@
Appendix G Proof of Proposition 5.4: Circulant Gaussian noise
The proof strategy for Proposition 5.4 closely follows that of the i.i.d. Gaussian case (Theorem 4.1), with appropriate modifications to handle circulant noise. The necessary assumptions and notations are introduced in Appendix G.1. Appendix G.2 establishes the asymptotic convergence of the EfN estimator as $M\to\infty$ under circulant Gaussian noise statistics. In Appendix G.3, we show that conditioning the EfN process on a single Fourier noise coefficient results in a cyclo-stationary process with a cosine trend. Appendix G.4 extends the vanishing imaginary part result from Appendix B.1 to the setting of circulant Gaussian noise. Similarly, Appendix G.5 extends the result of Appendix B.2, showing that the real part remains strictly positive in the circulant case. Finally, Appendix G.6 combines the results of the preceding sections to complete the proof of Proposition 5.4.
G.1 Preliminaries
@@
    Fourier phases distribution. For any $k$ such that $\lambda_{k}>0$, the Fourier coefficient $\mathsf{Y}_{i}[k]$ is a zero-mean, circularly symmetric complex Gaussian random variable. Hence, the phases $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d/2}$ are i.i.d. and uniformly distributed on $\pi,\pi)$ and independent of the magnitude $\{|\mathsf{Y}_{i}[k]|\}_{k=0}^{d/2}$.
@@
    $\mathsf{Y}_{i}[d-k]=\overline{\mathsf{Y}_{i}[k]},\quad\text{for }1\leq k\leq d-1.$ (G.1)
@@
G.2 The convergence of the Einstein from Noise estimator
@@
$\hat{\mathsf{X}}[k]$ $=\frac{1}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{\mathsf{Y}_{i}}[k]}\,e^{j\frac{2\pi k}{d}\mathsf{\hat{R}}_{i}}$ (G.2)
$=\frac{e^{j\phi_{\mathsf{X}}[k]}}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{\mathsf{Y}_{i}}[k]}\,e^{j\frac{2\pi k}{d}\mathsf{\hat{R}}_{i}}\,e^{-j\phi_{\mathsf{X}}[k]}$ (G.3)
$=\frac{e^{j\phi_{\mathsf{X}}[k]}}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{e,i}[k]},$ (G.4)
@@
$\mathsf{\hat{R}}_{i}$ $\triangleq\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}{\,\langle{y_{i}},\mathcal{T}_{r}x\rangle}$ (G.5)
$=\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}\ {\langle\mathcal{F}\left\{y_{i}\right\},{\mathcal{F}\left\{\mathcal{T}_{r}x\right\}}\rangle}$ (G.6)
$=\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}\sum_{k=0}^{d-1}{\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right)}.$ (G.7)
@@
$\phi_{e,i}[k]\triangleq\frac{2\pi k\mathsf{\hat{R}}_{i}}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k].$ (G.8)
@@
$\mathsf{S}_{i}[r]\triangleq\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (G.9)
so that $\mathsf{\hat{R}}_{i}=\operatorname*{arg\,max}_{0\leq r\leq d-1}\mathsf{S}_{i}[r]$. We note that for any $0\leq i\leq M-1$, the random vector $\mathsf{S}_{i}\triangleq(\mathsf{S}_{i}[0],\mathsf{S}_{i}[1],\ldots,\mathsf{S}_{i}[d-1])^{\top}$ is jointly Gaussian with zero mean and a circulant covariance matrix (as it is a Fourier transform of the convolution between $y_{i}$ and the template $x$). Hence, $\mathsf{S}_{i}$ forms a cyclo-stationary process. Applying the strong law of large numbers (SLLN), as $M\to\infty$, we have,
$\mathsf{\hat{X}}[k]\,e^{-j\phi_{\mathsf{X}}[k]}$ $=\frac{1}{M}\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\,e^{j\phi_{e,i}[k]}$ (G.10)
$\xrightarrow[]{\text{a.s.}}\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\,\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right],$ (G.11)
@@
$\mu_{\mathsf{A},k}$ $\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right],$ (G.12)
$\mu_{\mathsf{B},k}$ $\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right],$ (G.13)
@@
G.3 Conditioning on the Fourier frequency noise component
@@
Lemma G.2.
Let $\mathsf{S}_{i}$ be defined as in (G.9), and denote $\mathbb{E}\left[|\mathsf{Y}_{i}[k]|^{2}\right]=\lambda_{k}>0$ for each $k\in\{0,1,\ldots,d-1\}$. Then, for every $k\in\left\{1,2,\ldots,\tfrac{d}{2}-1,\tfrac{d}{2}+1,\ldots,d-1\right\}$, the random vector $\mathsf{S}_{i}$ conditioned on $\mathsf{Y}_{i}[k]$ is Gaussian:
$\mathsf{S}_{i}|\mathsf{Y}_{i}[k]\sim\mathcal{N}(\mathsf{\mu}_{k,i},\mathsf{\Sigma}_{k,i}),$ (G.14)
@@
$\mu_{k,i}[r]$ $\triangleq\mathbb{E}\left[\mathsf{S}_{i}[r]|\mathsf{Y}_{i}[k]\right]=2\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (G.15)
for $0\leq r\leq d-1$, and
@@
$=\sum_{\ell=0}^{d-1}\frac{\lambda_{\ell}}{2}\cdot|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$ (G.16)
for $0\leq r,s\leq d-1$, where $\widetilde{\mathsf{X}}_{k}$ is defined by:
$\widetilde{\mathsf{X}}_{k}[\ell]\triangleq\begin{cases}0&\mathsf{if}\ell=k,d-k,\\ \mathsf{X}[\ell]&\mathsf{if}\ell=0,d/2,\\ \sqrt{2}\cdot\mathsf{X}[\ell]&\mathsf{otherwise}.\end{cases}$ (G.17)
Note that the conditional process $\mathsf{S}_{i}|\mathsf{Y}_{i}[k]$ is Gaussian because it is given by a linear transform of i.i.d. Gaussian variables. Also, since its covariance matrix is circulant and depends only on the difference between the two indices, i.e., $\mathsf{\Sigma}_{k,i}[r,s]=\sigma_{k,i}[\left|r-s\right|]$, it is cycle-stationary with a cosine trend. The eigenvalues of this circulant matrix are given by the DFT of its first row, and thus its $\ell$-th eigenvalue equals $\lambda_{\ell}\cdot|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}$, for $0\leq\ell\leq d-1$.
Remark G.3.
@@
Proof of Lemma G.2.
We recall that if $\left\{y_{i}\right\}_{i=0}^{M-1}\sim\mathcal{N}(0,\Sigma)$, for symmetric circulant matrix $\Sigma$, then their DFT coefficients satisfy ${\{|\mathsf{Y}_{i}[k]|}\}_{k=0}^{d/2}$, and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d/2}$ are independent and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}\sim\mathsf{Unif}[-\pi,\pi)$. By definition of $\mathsf{S}_{i}$ (G.9), we have for every $k\neq 0,d/2$,
@@
$+\sum_{\ell\neq k,d-k}{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)},$ (G.18)
where we have used the property of $\mathsf{X}[k]=\overline{\mathsf{X}[d-k]},\ \mathsf{Y}_{i}[k]=\overline{\mathsf{Y}_{i}[d-k]}$. Clearly, as $\mathbb{E}\left[\mathsf{Y}_{i}\left[\ell\right]\right]=0$, for every $0\leq\ell\leq d-1$, we have,
$\mathbb{E}\left[{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)}\right]=0,$ (G.19)
for every $0\leq\ell\leq d-1$. Combining (G.18) and (G.19) results,
$\mu_{k,i}[r]$ $=\mathbb{E}\left[\mathsf{S}_{i}[r]|\mathsf{Y}_{i}[k]\right]=2\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right),$ (G.20)
@@
$\mathsf{\Sigma}_{k,i}[r,s]$ $\triangleq\mathbb{E}\left[\left(\mathsf{S}_{i}[r]-\mathbb{E}\mathsf{S}_{i}[r]\right)\left(\mathsf{S}_{i}[s]-\mathbb{E}\mathsf{S}_{i}[s]\right)|\mathsf{Y}_{i}[k]\right].$ (G.21)
@@
$=\sum_{\ell\neq k,d-k}{\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)},$ (G.22)
@@
$\mathcal{I}=\left\{1,2,\dots k-1,k+1,\dots,d/2-1\right\},$ (G.23)
@@
$+2\cdot\sum_{\ell\in\mathcal{I}}\left|\mathsf{X}[\ell]\right|\left|\mathsf{Y}_{i}[\ell]\right|\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right),$ (G.24)
where each one of the terms in the sum is independent. Since the terms in the sum on the r.h.s. of (G.24) are independent (i.e., $\mathbb{E}\left[\mathsf{Y}_{i}\left[\ell_{1}\right]\overline{\mathsf{Y}_{i}\left[\ell_{2}\right]}\right]=\mathbb{E}\left[\left|\mathsf{Y}_{i}\left[\ell_{1}\right]\right|^{2}\right]\delta_{\ell_{1},\ell_{2}}$), it follows that,
@@
$+4\cdot\mathbb{E}\left[{\sum_{\ell\in\mathcal{I}}{\left|\mathsf{X}[\ell]\right|^{2}\left|\mathsf{Y}_{i}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell r}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\cos\left(\frac{2\pi\ell s}{d}+\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)}}\right].$ (G.25)
The expectation value in (G.25) is composed of the multiplications of cosines. Applying trigonometric identities, we obtain,
@@
$=\frac{1}{2}\cos\left(\frac{2\pi\ell(r-s)}{d}\right)+\frac{1}{2}\cos\left(\frac{2\pi\ell(r+s)}{d}+2\left(\phi_{\mathsf{Y}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\right),$ (G.26)
for every $0\leq r,s\leq d-1$. Now, since the sequences $\left\{|\mathsf{Y}_{i}[\ell]|\right\}_{\ell=0}^{d/2}$ and $\left\{\phi_{\mathsf{Y}_{i}}[\ell]\right\}_{\ell=0}^{d/2}$ are independent random variables, with $\mathbb{E}\left[\left|\mathsf{Y}_{i}[k]\right|^{2}\right]=\lambda_{k}$ and phases $\phi_{\mathsf{Y}_{i}}[k]$ uniformly distributed over $\pi,\pi)$, and by applying the trigonometric identity (G.26), it follows that,